I. Determine whether it is categorical, quantitative discrete, or quantitative continuous.

a.      Technology ownership: Do you own a computer?

b.     Time (per week in minutes) spent using a computer for writing documents (word processing).

c.      Which social networking site(s) are you a member? (Facebook, Myspace, Friendster, etc.)

II. True or False

1. If the time to wait for pharmacy help has a uniform distribution from 0 minutes to 30 minutes, then 33% of the customers are expected to wait more than 20 minutes.

2.      If X has a Binomial (50, 0.7) distribution, then the criteria to use the normal approximation are met.


1
Expert's answer
2021-07-16T14:31:25-0400

I.

(a) categorical

Either you have a computer, or you do not.


(b) quantitative continuous

Time is continuous.


(c) categorical

There can be multiple categories along with given three(Facebook, Myspace, Friendster)



II.

(1) For uniform distribution

"Pr(X\leq x)=\dfrac{x-a}{b-1}"


Therefore,

"Pr(X>x)= 1-Pr(X\leq x)"

"= 1-\dfrac{20-0}{30-0}\\=1-\dfrac{2}{3}\\=\dfrac{1}{3}=33\%"


Hence, the correct choice is TRUE


(2) Criteria to use normal approximation to binomial

To use normal approximation:

"n*p\geq5\ \ and\ \ n*q\geq 5"


Thus;

"50*0.7 =35\\50*(1-0.7)= 15"


Hence, the conditions are met, the correct choice is TRUE